Market price of bread before tax =$100 per unit
Unit tax =$30
Market price of bread after tax came into effect = $110
Quantity of bread sold after tax came into effect=1000
1. Calculate the sellers new after tax price
2. Calculate government s total tax revenue
3. Calculate tax incidence borne by buyers
4. Calculate tax incidence borne by sellers
1
Expert's answer
2019-04-25T09:59:47-0400

1. The sellers new after tax price is:

Ps = 110 - 30 = $80.

2. Government's total tax revenue is:

TR = $30×1000 = $30000.

3. Tax incidence borne by buyers is:

(110 - 100)×1000 = $10000.

4. Tax incidence borne by sellers is:

(100 - 80)×1000 = $20000.